OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 5. Using complete sentences, explain the procedure used to solve this equation

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

using the log laws \[\log(ab) = \log(a) + \log(b)\] and \[\log _{a} a = 1\] so look at \[\log _{16} (32) = \log _{16}(2\times16) = \log _{16}(2) + 1\] and \[2 = 16^{1/4}\] so \[\log _{16}(16^{\frac{1}4}) = 1/4\] so the problem is now \[\frac{1}{4} + 1 = x + 5\] or \[\frac{5}{4} = x + 5\] solve for x

OpenStudy (anonymous):

thank you so much